The sky is clear.

Summary

There are total ~2.8 hours scientific exposure. 32 min J0900+2234 J H K image 25 min J0900+2234 H+K spectra (bad seeing) 30 min 3C220 J H K image 42 min cosmos4 H+K spectra 21 min cosmos3 H+K spectra (band seeing+ thin cloud) 15 min IMACS1132 J H K image

23:30
Refreshing the detector, and the counts at the right corner is higher (~100 ADU) (due to the lights in the dome)

00:30
#03-12 flat K band 2s ~13000 counts

00:47
#13-22 flat H band 4s ~15000 counts.
#23-32 flat J band 7s ~10000 counts.

01:05
Check the pointing, collimate the telescope the seeing looks not good 1.8, collimate with a star around the J2329-0301, seeing 1.5.

01:40
Go to J2329-0301. When do the acquisition imaging, the images are not sent the skycat automatically (this happened serveral times last night, even after restarted skycat)

02:08
Start scientific exporesure of J2329-0301, seeing is 1.0 #52 is good

02:22
Marco restart the data acquiration system, and we begin to lose our data from #53, the second exposure. #53 is bad

02:30
Go back to acquiration again frown, sad smile

02:38
Go to the J2329-0301. MOS section red. Error shows up, when trying to put the 1" mask into the beam Marco waked up Walter. Nancy is working on the problem.

03:30
The problem is solved. J2329-0301 is too low.

03:32
David goes up to set the rotator manually

04:00
Do collimation. Meet an error when put the mask out of the beam, ignore it, initialize.
Can not find a star for guiding. Check the pointing. The pointing is good. 04:00 Do collimation. Meet an error when put the mask out of the beam, ignore it, initialize. Can not find a star for guiding. Check the pointing. The pointing is good. David reset the AGW. Probably AGW issues. The persion standard star catalog has wrong the guide star coordination (BS9105_GS12). And the CCD seems dead. Power recycle the AGW. Go to another persion star. It works now.

04:56
We go to the acquisition of J0203+0012

05:13
We lose the guide star. Seeing 0.7

05:21
Go back to J0203+0012, can not get the guide star. Go back to the persion standard star, and find the guide star. Try another guide star, it is too faint.

05:33
Still can not find the guide star. Do pointing

05:38
Now we are happy. and unhappy again, because the probe lose the guide star again.

05:54
We go to J band image of J0900+2234, seeing is 0.5" # 87-88 The guide star is our of range during the dithering after #88 Change the script for a small offset value. #90-99 J image

06:38
go to H band imaging of J0900+2234 Can not find the guide then we go to check pointing 20" difference in AC direction. (Only expected in a large temperature jump)

06:49
#103-112 H band image of J0900+2234

07:11
#113-122 Ks band image of J0900+2234 Compare with the H band image, we can see the telescope drift a little bit.

07:33
#128-129 Spectra of J0900+2234 (150*2). please not the object name in the head is wrong. The source is not in the slit, 2 exposures taken. Everything is the same as last night, could be due to the bad pointing? Then we take a 2 minutes exposure and can see the target and put them in the slit. #133-137 Can not see anything in the slit and seeing is very bad, 1.0-2.0", should redo this exposure

08:53
We go to the J H Ks imaging of 3C220

#138-147 J band imaging #148-157 H band imaging #158-167 Ks band imaging

09:48
Go to the cosmos4, seeing 1" #174-187 H+K spectra of cosmos4

11:21
go to the standard star 69 Leo. #188-190,

lamp #191-196 ,

flats #200,

11:45
Go to the cosmos3, seeing 1" Guide star is too faint, the guider can not find. Try another brighter one, can find the star but the star image can not converge Go to nearby persion standard star to collimate. We waiting for the timeout of the guider to collimate. Ask David to stop the guider, still can not commit the telescope. The guide star begins to converge, seeing 1".

12:15
#206-212 (bad condition) science exposure of cosmos3 starts. The seeing getting lousy around 12:30. and cloud pass by around 12:45 Should redo this exposure

12:49
Stop the exposure due to the bad codition. Go to J H K imaging of MACS1132 seeing is 1" #214-223 J band imaging #224-229 H band imaging

13:22
lost the guide star, totally cloudy. We are done.

The light in the dorm is on, when we take the dark, since the daycrew needs to work on the HBS system. The dark before 284 may be bad. .

-- FuyanBian - 06 Jan 2010
Topic revision: r1 - 07 Jan 2010, FuyanBian
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback